OpenStudy (anonymous):

evaluate determinants. top row -4,5,15 2nd row 8,-7,16 3rd row -4,21,-2

OpenStudy (unklerhaukus):

\[ det\left[\begin{matrix}-4&5 &15 \\ 8&-7&16\\ -4&21 & -2\end{matrix}\right] \] \[=-4((-7\times-2)-(16 \times21))-5((8\times-2)-(-4\times 16))+15((8\times21)-(-7 \times -4)\] \[=3148\]

OpenStudy (anonymous):

Use expansion by minors along the top row, for instance. Don't forget to flip the sign on the second term.\[-4\begin{vmatrix}-7 & 16 \\ 21 & -2\end{vmatrix}-5\begin{vmatrix}-8 & 16 \\ -4 & -2\end{vmatrix}+15\begin{vmatrix}8 & -7 \\ -4 & 21\end{vmatrix}\]Now, you can evaluate the smaller determinants by the following formula, which you definitely should make an effort to be second nature.\[\begin{vmatrix}a & b \\ c & d\end{vmatrix} = ad-bc\]